# Data-Driven LQR with Finite-Time Experiments via Extremum-Seeking Policy Iteration

### v2

@@ VII Conclusions
where $(a)$ uses the Taylor expansion of $J(\cdot)$ about $K_{\textsc{av}}$ evaluated at $K-\gamma G(K_{\textsc{av}})+\gamma p^{\delta,T}(K_{\textsc{av}})$, (45), (47b), and the Cauchy-Schwarz inequality, while $(b)$ rearranges the terms and uses $\frac{1}{2}\left\|G(K_{\textsc{av}})-p^{\delta,T}(K_{\textsc{av}})\right\|^{2}\leq\left\|G(K_{\textsc{av}})\right\|^{2}+\left\|p^{\delta,T}(K_{\textsc{av}})\right\|^{2}$. Let us arbitrarily fix $\eta\in(0,1)$ and define $\bar{\gamma}_{2}:=\min\{\bar{\gamma}_{1},\frac{1-\eta}{\beta_{2}}\}$. Then, for all $\gamma\in(0,\bar{\gamma}_{2})$, we can bound (48) as
$\Delta V_{K}(K_{\textsc{av}})$ $\leq-\gamma\eta\left\|G(K_{\textsc{av}})\right\|^{2}+\gamma\left\|G(K_{\textsc{av}})\right\|\left\|p^{\delta,T}(K_{\textsc{av}})\right\|$
@@
where in $(a)$ we use the results (41) and (43) to bound $G(K_{\textsc{av}})$ and $\left\|p^{\delta,T}(K_{\textsc{av}})\right\|$ over the compact set $\Omega^{K}_{c_{0}}$. Now, in order to handle also the dynamics (16a), let us introduce $\tilde{\mathrm{z}}_{\textsc{av}}\in\mathbb{R}$ defined as
@@
$g^{\delta}(K):=-J_{\textsc{av}}^{\delta}(K_{\textsc{av}}-\gamma G(K_{\textsc{av}})+\gamma p^{\delta,T}(K_{\textsc{av}}^{k}))+J_{\textsc{av}}^{\delta}(K_{\textsc{av}}).$ (52)
@@
$V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})=\frac{1}{2}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}.$ (53)
Hence, by evaluating the increment $\Delta V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}}):=V_{\mathrm{z}}((1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}+g^{\delta}(K_{\textsc{av}}))-V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})$ of $V_{\mathrm{z}}$ along the trajectories of (51a), it holds
$\Delta V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})$ $=-\gamma\left(1-\frac{\gamma}{2}\right)\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+(1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}^{k}g^{\delta}(K_{\textsc{av}})$
$\hskip 11.38092pt+\frac{1}{2}\left\|g^{\delta}(K_{\textsc{av}})\right\|^{2}.$ (54)
@@
$\min_{K_{\textsc{av}}\in\Omega^{K}_{c_{0}}}\left\|\nabla g^{\delta}(K_{\textsc{av}})\right\|\leq\beta_{J_{\textsc{av}}},$ (55)
@@
$\Delta V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})$ $\leq-\gamma\left(1-\frac{\gamma}{2}\right)\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}-\gamma\left(1-\frac{\gamma}{2}\right)\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}$
$\hskip 11.38092pt+(1-\gamma)\gamma\beta_{J_{\textsc{av}}}\left\|\tilde{\mathrm{z}}_{\textsc{av}}^{k}\right\|\left\|G(K_{\textsc{av}})\right\|$
@@
$\hskip 11.38092pt+\delta^{2}(1-\gamma)\gamma\beta_{J_{\textsc{av}}}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|\beta_{e}$
$\hskip 11.38092pt+\delta^{4}\gamma^{2}\beta_{J_{\textsc{av}}}^{2}\beta_{e}^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-\gamma\left(1-\frac{\gamma(1+\kappa)}{2}\right)\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}$
@@
$\hskip 11.38092pt+\delta^{4}\left(\gamma^{2}+\frac{1}{2\kappa}(1-\gamma)^{2}\right)\beta_{J_{\textsc{av}}}^{2}\beta_{e}^{2},$ (57)
where in $(a)$ we use (43) to bound $\left\|p^{\delta,T}(K_{\textsc{av}})\right\|$, while in $(b)$ we use the Young’s inequality with an arbitrarily fixed parameter $\kappa>0$ to handle the term $\delta^{2}(1-\gamma)\gamma\beta_{J_{\textsc{av}}}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|\beta_{e}$. Now, let us introduce a function to compactly contain all the terms due to the approximation error $p^{\delta,T}$, namely, let us define
$\tilde{V}(\gamma,\delta,\lambda):=\delta^{2}\gamma\beta_{0}\beta_{p}^{2}+\delta^{4}\Big{(}\gamma^{2}\beta_{2}\beta_{p}^{2}+\tfrac{1}{\lambda}\Big{(}\gamma^{2}+\frac{1}{2\kappa}\beta_{J_{\textsc{av}}}^{2}\beta_{p}^{2}\Big{)}\Big{)}.$ (58)
Then, let us consider the function $V_{\lambda}$ (cf. (18)) and evaluate its increment $\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}):=V_{\lambda}((1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}+g^{\delta}(K_{\textsc{av}}),K_{\textsc{av}}-\gamma G(K_{\textsc{av}})+\gamma p^{\delta,T}(K_{\textsc{av}}))-V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ along the trajectories of (51). By using the bounds (49)-(57) and the definition of $\tilde{V}$ (cf. (58)), we get
@@
$\leq-\gamma\begin{bmatrix}\tilde{\mathrm{z}}_{\textsc{av}}\\ G(K_{\textsc{av}})\end{bmatrix}^{\top}\mathcal{Q}(\gamma,\lambda)\begin{bmatrix}\tilde{\mathrm{z}}_{\textsc{av}}\\ G(K_{\textsc{av}})\end{bmatrix}+\tilde{V}(\gamma,\delta,\lambda),$ (59)
@@
$\mathcal{Q}(\gamma,\lambda):=\begin{bmatrix}\frac{1}{\lambda}\left(1-\frac{\gamma(1+\kappa)}{2}\right)&-\frac{(1-\gamma)\beta_{J_{\textsc{av}}}}{2\lambda}\\ -\frac{(1-\gamma)\beta_{J_{\textsc{av}}}}{2\lambda}&\eta-\gamma\frac{\beta_{J_{\textsc{av}}}^{2}}{2\lambda}\end{bmatrix}.$
Let us impose the positive definiteness of the top-left entry of $\mathcal{Q}(\gamma,\lambda)$. To this end, let us arbitrarily fix $\nu\in(0,2)$ and define $\bar{\gamma}_{1}:=\min\{\bar{\gamma}_{0},\bar{\gamma}_{2},(2-\nu)/(1+\kappa)\}$. Then, by Sylvester Criterion, for all $\gamma\in(0,\bar{\gamma}_{1})$, it holds
$\mathcal{Q}(\gamma,\lambda)\succeq\begin{bmatrix}\frac{\nu}{2\lambda}&-\frac{(1-\gamma)\beta_{J_{\textsc{av}}}^{2}}{2\lambda}\\ -\frac{(1-\gamma)\beta_{J_{\textsc{av}}}}{2\lambda}&\eta-\frac{\beta_{J_{\textsc{av}}}^{2}}{2\lambda}\end{bmatrix}.$
Now, let us impose the positive definiteness of the entire matrix $\mathcal{Q}(\gamma,\lambda)$. To this end, let us arbitrarily fix $\tilde{\eta}\in(0,\eta)$ and $\tilde{\nu}\in(0,\nu)$ and define
$\bar{\lambda}:=\max\left\{\frac{\beta_{J_{\textsc{av}}}^{2}+2\beta_{J_{\textsc{av}}}(\nu-\tilde{\nu})}{4(\eta-\tilde{\eta})(\nu-\tilde{\nu})},1\right\}.$ (60)
Then, we arbitrarily fix $\lambda\geq\bar{\lambda}$ and, by Sylvester Criterion, it holds
@@
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma\left(\frac{\tilde{\nu}}{2\lambda}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+\tilde{\eta}\left\|G(K_{\textsc{av}})\right\|^{2}\right)$
@@
Now, by applying the gradient dominance property of $J$ (cf. (46)), we note that
$-\frac{\tilde{\nu}}{2\lambda}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}-\tilde{\eta}\left\|G(K_{\textsc{av}})\right\|^{2}$
$\leq-\frac{\tilde{\nu}}{2\lambda}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}-\frac{\tilde{\eta}}{\mu}\left(J(K_{\textsc{av}})-J(K^{\star})\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}-vV_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}),$ (62)
where in $(a)$ we use the definition of $V_{\lambda}$ (cf. (18)) and $v:=\max\{\tilde{\nu},\tilde{\eta}/\mu\}$. Then, by using (62), we can further bound (61) according to
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\leq-\gamma vV_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})+\tilde{V}(\bar{\gamma}_{1},\delta,\lambda).$ (63)
Now, without loss of generality, we assume $\rho_{\textsc{av}}\leq c_{0}$. Indeed, one may always recover such a condition by using $\max\{\rho_{\textsc{av}},c_{0}\}$ in place of $c_{0}$. Subsequently, given $\tilde{v}\in(0,v)$, let us introduce $\bar{\delta}_{3}>0$ defined as
$\bar{\delta}_{3}:=\min\left\{\sqrt{\frac{\left(v-\tilde{v}\right)\rho_{\textsc{av}}}{\bar{\gamma}_{1}\beta_{0}\beta_{p}+\bar{\gamma}_{1}^{2}\beta_{2}\beta_{p}^{2}+\tfrac{1}{\lambda}\Big{(}\bar{\gamma}_{1}^{2}+\frac{1}{2\kappa}\beta_{J_{\textsc{av}}}^{2}\beta_{p}^{2}\Big{)}}},1\right\}.$
@@
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma\tilde{v}V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}),$ (64)
for all $(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\in\mathbb{R}\times\mathcal{K}$ such that $V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\geq\rho_{\textsc{av}}$. Although (64) seems to conclude the proof, we recall that it has been obtained by assuming $K_{\textsc{av}}\in\tilde{\Omega}_{c_{0}}$. In other words, since $(\tilde{\mathrm{z}}_{\textsc{av}}^{0},K_{\textsc{av}}^{0})\in\tilde{\Omega}_{c_{0}}$ by definition of $c_{0}$, to conclude the proof we only need to prove that the set ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ is forward-invariant for system (51). To this end, consider $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\tilde{\Omega}_{c_{0}}$ and, in light of the definition of $V_{\lambda}$ (cf. (18)), we note that
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{1}{2}\left\|\tilde{\mathrm{z}}_{\textsc{av}}^{k}\right\|^{2}+J(K_{\textsc{av}}^{k})-J(K^{\star}),$ (65)
where in $(a)$ we use the fact that the right-hand side of (64) is non-positive for all $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\mathbb{R}\times\mathcal{K}$ such that $V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\geq\rho_{\textsc{av}}$, while in $(b)$ we use the fact that $\lambda\geq 1$. By looking at the definition of $\tilde{\Omega}$ in (40), the inequality (65) proves the desired invariance property of ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ and the proof concludes.

### v3

@@ VII Conclusions
where $(a)$ uses the Taylor expansion of $J(\cdot)$ about $K_{\textsc{av}}$ evaluated at $K-\gamma G(K_{\textsc{av}})+\gamma p^{\delta,T}(K_{\textsc{av}})$, (46), (48), and the Cauchy-Schwarz inequality, while $(b)$ rearranges the terms and uses $\frac{1}{2}\left\|G(K_{\textsc{av}})-p^{\delta,T}(K_{\textsc{av}})\right\|^{2}\leq\left\|G(K_{\textsc{av}})\right\|^{2}+\left\|p^{\delta,T}(K_{\textsc{av}})\right\|^{2}$. Let us arbitrarily fix $\eta\in(0,1)$ and define $\bar{\gamma}_{2}:=\min\{\bar{\gamma}_{1},\frac{1-\eta}{\beta_{2}}\}$. Then, for all $\gamma\in(0,\bar{\gamma}_{2})$, we can bound (49) as
$\Delta V_{K_{\textsc{av}}}(K_{\textsc{av}})$ $\leq-\gamma\eta\left\|G(K_{\textsc{av}})\right\|^{2}+\gamma\left\|G(K_{\textsc{av}})\right\|\left\|p^{\delta,T}(K_{\textsc{av}})\right\|$
@@
where in $(a)$ we use (42) and (44) to bound $G(K_{\textsc{av}})$ and $\left\|p^{\delta,T}(K_{\textsc{av}})\right\|$ over the compact set $\Omega^{K}_{c_{0}}$. Now, in order to handle also the dynamics (17a), let us introduce $\tilde{\mathrm{z}}_{\textsc{av}}\in\mathbb{R}$ defined as
@@
$g^{\delta}(K):=J_{\textsc{av}}^{\delta}(K)-J_{\textsc{av}}^{\delta}(K-\gamma G(K)+\gamma p^{\delta,T}(K)).$ (53)
@@
$V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})=\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}/2.$ (54)
Hence, the increment $\Delta V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}}):=V_{\mathrm{z}}((1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}+g^{\delta}(K_{\textsc{av}}))-V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})$ of $V_{\mathrm{z}}$ along the trajectories of (52a) reads as
$\Delta V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})$ $=-\gamma(1-\gamma/2)\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+(1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}^{k}g^{\delta}(K_{\textsc{av}})$
$\hskip 11.38092pt+1/2\left\|g^{\delta}(K_{\textsc{av}})\right\|^{2}.$ (55)
@@
$\min_{K_{\textsc{av}}\in\tilde{\Omega}_{c_{0}}}\left\|\nabla J_{\textsc{av}}^{\delta}(K_{\textsc{av}})\right\|\leq\beta_{J_{\textsc{av}}},$ (56)
@@
$\Delta V_{\mathrm{z}}(\tilde{\mathrm{z}}_{\textsc{av}})$ $\leq-\gamma\big{(}1-\tfrac{\gamma}{2}\big{)}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+(1-\gamma)\gamma\beta_{J_{\textsc{av}}}\!\!\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|\!\left\|G(K_{\textsc{av}})\right\|$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\!\!-\gamma\big{(}1-\tfrac{\gamma}{2}\big{)}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+(1-\gamma)\gamma\beta_{J_{\textsc{av}}}\!\!\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|\!\left\|G(K_{\textsc{av}})\right\|$
@@
$\hskip 11.38092pt+\delta^{2}(1-\gamma)\gamma\beta_{J_{\textsc{av}}}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|\beta_{e}+\delta^{4}\gamma^{2}\beta_{J_{\textsc{av}}}^{2}\beta_{e}^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-\gamma(1-\kappa/2-\gamma/2)\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}$
@@
$\hskip 11.38092pt+\delta^{4}\left(\gamma^{2}+\gamma/(2\kappa)(1-\gamma)^{2}\right)\beta_{J_{\textsc{av}}}^{2}\beta_{e}^{2},$ (58)
where in $(a)$ we use (44) to bound $\left\|p^{\delta,T}(K_{\textsc{av}})\right\|$, while in $(b)$ we use the Young’s inequality with an arbitrarily fixed parameter $\kappa\in(0,2)$ and write $\delta^{2}(1-\gamma)\gamma\beta_{J_{\textsc{av}}}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|\beta_{e}\leq\gamma(\frac{\kappa}{2}\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+\frac{1}{2\kappa}\delta^{4}(1-\gamma)^{2}\beta_{J_{\textsc{av}}}^{2}\beta_{e}^{2})$. Now, let us compactly write all the terms due to the approximation error $p^{\delta,T}$ by introducing $\tilde{V}:\mathbb{R}\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ defined as
$\tilde{V}(\gamma,\delta,\lambda)\!:=\!\delta^{2}\beta_{0}\beta_{p}^{2}\!+\!\delta^{4}(\gamma\beta_{2}\beta_{p}^{2}\!+\!\tfrac{1}{\lambda}(\gamma\!+\!\tfrac{1}{2\kappa}\beta_{J_{\textsc{av}}}^{2}\beta_{p}^{2})).\!\!$ (59)
Then, let us $V_{\lambda}$ (cf. (19)) and evaluate its increment $\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}):=V_{\lambda}((1-\gamma)\tilde{\mathrm{z}}_{\textsc{av}}+g^{\delta}(K_{\textsc{av}}),K_{\textsc{av}}-\gamma G(K_{\textsc{av}})+\gamma p^{\delta,T}(K_{\textsc{av}}))-V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ along the trajectories of (52). By using (50), (58), and the definition of $\tilde{V}$ (cf. (59)), we get
@@
$\leq-\gamma\begin{bmatrix}\tilde{\mathrm{z}}_{\textsc{av}}\\ G(K_{\textsc{av}})\end{bmatrix}^{\top}\mathcal{Q}(\gamma,\lambda)\begin{bmatrix}\tilde{\mathrm{z}}_{\textsc{av}}\\ G(K_{\textsc{av}})\end{bmatrix}+\gamma\tilde{V}(\gamma,\delta,\lambda),$ (60)
@@
$\mathcal{Q}(\gamma,\lambda):=\begin{bmatrix}\frac{1}{\lambda}\big{(}1-\frac{\kappa}{2}-\frac{\gamma}{2}\big{)}&-\frac{(1-\gamma)\beta_{J_{\textsc{av}}}}{2\lambda}\\ -\frac{(1-\gamma)\beta_{J_{\textsc{av}}}}{2\lambda}&\eta-\gamma\frac{\beta_{J_{\textsc{av}}}^{2}}{\lambda}\end{bmatrix}.$
Let us impose the positive definiteness of the top-left entry of $\mathcal{Q}(\gamma,\lambda)$. To this end, let us arbitrarily fix $\nu\in(0,1-\kappa/2)$ and define $\bar{\gamma}_{1}:=\min\{\bar{\gamma}_{0},\bar{\gamma}_{2},2(1-\kappa/2-\nu)\}$. Then, by Sylvester Criterion, for all $\gamma\in(0,\bar{\gamma}_{1})$, it holds
$\mathcal{Q}(\gamma,\lambda)\succeq\begin{bmatrix}\frac{\nu}{\lambda}&-\frac{(1-\gamma)\beta_{J_{\textsc{av}}}}{2\lambda}\\ -\frac{(1-\gamma)\beta_{J_{\textsc{av}}}}{2\lambda}&\eta-\gamma\frac{\beta_{J_{\textsc{av}}}^{2}}{\lambda}\end{bmatrix}.$
Now, let us impose the positive definiteness of $\mathcal{Q}(\gamma,\lambda)$. To this end, we arbitrarily fix $\tilde{\eta}\in(0,\eta)$, $\tilde{\nu}\in(0,\nu)$, and define
$\bar{\lambda}:=\max\left\{\frac{(1+\bar{\gamma}_{1}^{2})\beta_{J_{\textsc{av}}}^{2}+4\bar{\gamma}_{1}\beta_{J_{\textsc{av}}}^{2}(\nu-\tilde{\nu})}{4(\eta-\tilde{\eta})(\nu-\tilde{\nu})},1\right\}.$ (61)
Then, we arbitrarily fix $\lambda\geq\bar{\lambda}$ and the Sylvester Criterion yelds
@@
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma(\tilde{\nu}/\lambda\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}+\tilde{\eta}\left\|G(K_{\textsc{av}})\right\|^{2})$
@@
The gradient dominance property of $J$ (cf. (47)) leads to
$-\tilde{\nu}/\lambda\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}-\tilde{\eta}\left\|G(K_{\textsc{av}})\right\|^{2}$
$\leq-\tilde{\nu}/\lambda\left\|\tilde{\mathrm{z}}_{\textsc{av}}\right\|^{2}-\tilde{\eta}/\mu\left(J(K_{\textsc{av}})-J(K^{\star})\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}-aV(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}}),$ (63)
where in $(a)$ we use the definition of $V_{\lambda}$ (cf. (19)) and $a:=\min\{2\tilde{\nu},\tilde{\eta}/\mu\}$. Then, by using (63), we further bound (62) as
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})\leq-\gamma aV_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})+\gamma\tilde{V}(\gamma,\delta,\lambda).$ (64)
Now, without loss of generality, we assume $\rho_{\textsc{av}}\leq c_{0}$. Indeed, one may always recover such a condition by using $\max\{\rho_{\textsc{av}},c_{0}\}$ in place of $c_{0}$. Then, let us define $\bar{\delta}_{3}>0$ as
$\bar{\delta}_{3}:=\min\left\{\sqrt{\frac{a\rho_{\textsc{av}}}{\delta^{2}\beta_{0}\beta_{p}^{2}\!+\!\delta^{4}(\bar{\gamma}_{1}\beta_{2}\beta_{p}^{2}\!+\!\tfrac{1}{\lambda}(\bar{\gamma}_{1}\!+\!\tfrac{1}{2\kappa}\beta_{J_{\textsc{av}}}^{2}\beta_{p}^{2}))}},1\right\}.$
@@
$\Delta V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})$ $\leq-\gamma aV_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}},K_{\textsc{av}})+\gamma a\rho_{\textsc{av}}.$ (65)
Although (65) seems to conclude the proof, we recall that it has been obtained by assuming $K_{\textsc{av}}\in\tilde{\Omega}_{c_{0}}$. In other words, since $(\tilde{\mathrm{z}}_{\textsc{av}}^{0},K_{\textsc{av}}^{0})\in\tilde{\Omega}_{c_{0}}$ by definition of $c_{0}$, to conclude the proof we only need to prove that the set ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ is forward-invariant for system (52). To this end, consider $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\tilde{\Omega}_{c_{0}}$ and, in light of the definition of $V_{\lambda}$ (cf. (19)), we note that
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})$ (66)
$\stackrel{{\scriptstyle(b)}}{{\leq}}1/2\left\|\tilde{\mathrm{z}}_{\textsc{av}}^{k}\right\|^{2}+J(K_{\textsc{av}}^{k})-J(K^{\star})\stackrel{{\scriptstyle(c)}}{{\leq}}c_{0},$
where in $(a)$ we use the fact that the right-hand side of (65) is non-positive for all $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\mathbb{R}\times\mathcal{K}$ such that $V_{\lambda}(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\geq\rho_{\textsc{av}}$, in $(b)$ we use the fact that $\lambda\geq 1$, while $(c)$ follows by the definition of $\tilde{\Omega}_{c_{0}}$ (cf. (41)) and that $(\tilde{\mathrm{z}}_{\textsc{av}}^{k},K_{\textsc{av}}^{k})\in\tilde{\Omega}_{c_{0}}$ by hypothesis. The inequality (66) proves the desired invariance property of ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ and the proof concludes.